# Imaginary eigenvalues of Hermitian Hamiltonian with an inverted potential well and transition to the real spectrum at exceptional point by a non-Hermitian interaction

### v1

@@ Abstract
February 9, 2024 February 9, 2024 February 9, 2024 February 9, 2024 February 9, 2024 February 9, 2024 February 9, 2024 February 9, 2024
@@ I Introduction
In quantum mechanics the Hamiltonian has to be Hermitian to guarantee real energy spectrum for the closed systems. The non-Hermitian Hamiltonian describes usually open systems, which exchange energy continuously with external source and thus possess complex energy spectrum. Bender and his collaborators proposed for the first time a creative finding that the Hermitian Hamiltonian is sufficient but not necessary to have real eigenvalues CM98 ; CM99 ; CM02 . According to them the non-Hermitian Hamiltonian with real spectrum has the parity-time ($PT$) reflection symmetry CM98 ; CM99 ; CM02 ; GuResult ; GuAnn ; LiuResult . Mostafazadeh named the non-Hermitian Hamiltonian with real spectrum as pseudo-Hermitian, which satisfies a necessary condition Mosta1 ; Mosta2 ; Mosta3 . While the $PT$ symmetry is neither necessary nor sufficient condition for a non-Hermitian Hamiltonian to have real spectrum. It is found that the $PT$asymmetric Hamiltonian can also possess real spectraLiuPysica ; Amaa .
For the $PT$symmetric non-Hermitian Hamiltonians there exists an exceptional point, which separates the unbroken $PT$ symmetry region with pure real-spectrum and a broken $PT$ symmetry region, where eigenvalues are complexCM98 ; CM04 ; CM07 . This transition at the exceptional point has been observed in numerous laboratory experimentsCM18 ; DN18 for various non-Hermitian Hamiltonians. The non-Hermitian Hamiltonian has become a rapidly developing subject in different branches of physicsCH13 ; ZZ16 ; SB12 ; CM13 ; CE10 ; YW19 . The spectrum property becomes complicated at the exceptional pointsLC14 ; PG15 ; HX16 ; WL17 due to the symmetry breaking, which can be modulated by dissipative mediaSK19 ; JZ18 ; YZ17 ; FQ18 . The exceptional point and complex domain of eigenvalues have been well studied particularly in relation with optical systemsCM18 ; DN18 .
@@ II Hermitian Hamiltonian with an inverted potential well and imaginary spectrum
in which the subscripts “$r$”, “$l$” denote respectively the “ket” and “bra” states. The two sets of eigenstates are mutually orthonormal
@@
for both the “bra” and “ket” states ($\gamma=l,r$). The state densities $\langle x|\widehat{\rho}|x\rangle$, $\langle x|\widehat{\rho}^{{\dagger}}|x\rangle$ are conserved quantities satisfying the basic requirement of quantum mechanics.
@@
with $\gamma=r,l$ respectively for the “ket” and “bra” states. Indeed the Hermitian Hamiltonian with imaginary eigenvalues Eq.(11) has a dual-set of eigenstates in agreement with the general theory Eq.(1).
@@
$|n\rangle_{l}$ $=$ $\frac{\left(\widehat{b}_{-}\right)^{n}}{i^{n}\sqrt{n!}}|0\rangle_{l}.$
@@
$\psi_{r,n}\left(x,\Omega\right)$ $=$ $\frac{1}{\pi^{\frac{1}{4}}\sqrt{2^{n}n!}}\sqrt{i^{\left(n+\frac{1}{2}\right)}}\left(x-i\frac{d}{dx}\right)^{n}e^{-i\frac{x^{2}}{2}},$
$\psi_{l,n}\left(x,\Omega\right)$ $=$ $\frac{1}{\pi^{\frac{1}{4}}\sqrt{2^{n}n!i^{\left(n-\frac{1}{2}\right)}}}\left(x+i\frac{d}{dx}\right)^{n}e^{i\frac{x^{2}}{2}},$ (12)
@@ III Transition from imaginary to real domains of spectrum and the exceptional point
where the real-value parameter $\eta$ is to be determined. The $SU(1,1)$ generators are realized by the imaginay-frequency boson operators in Eq.(5). It is easy to check that the transformation operator Eq.(17) is Hermitian $\widehat{R}=\widehat{R}^{{\dagger}}$. Using the transformation relations GuAnn ; LLM
$\widehat{R}\widehat{S}_{+}\widehat{R}^{-1}$ $=$ $\widehat{S}_{+}\cosh^{2}\frac{\eta}{2}-\widehat{S}_{-}\sinh^{2}\frac{\eta}{2}-i\widehat{S}_{z}\sinh\eta,$
$\widehat{R}\widehat{S}_{-}\widehat{R}^{-1}$ $=$ $\widehat{S}_{-}\cosh^{2}\frac{\eta}{2}-\widehat{S}_{+}\sinh^{2}\frac{\eta}{2}+i\widehat{S}_{z}\sinh\eta,$
$\widehat{R}\widehat{S}_{z}\widehat{R}^{-1}$ $=$ $\widehat{S}_{z}\cosh\eta+\frac{i}{2}(\widehat{S}_{+}-\widehat{S}_{-})\sinh\eta,$
@@
$\sinh\eta=\pm\frac{G}{\sqrt{\Omega^{2}-G^{2}}},$ (20)
@@
respectively for the “ket” and “bra” states, $\gamma=r,l$. In our case it is simply the eigenstate of imaginary-frequency boson-number $\widehat{n}_{I}$
@@
defined in equation Eq.(9). The eigenstates of the original non-Hermitian Hamiltonians in equations Eq.(13),Eq.(14) are obtained respectively by the inverse transforms
@@
Figure 2: The lowest layers ($n=0,1,2)$ of energy spectrum as functions of the interaction constant $G$. The imaginary eigenvalues below $G_{c}$ have two branches respectively for the “ket” and “bra” states and vanish at $G_{c}$. The spectrum becomes real beyond the exceptional point.
With the coordinate representation of $SU(1,1)$ generators $\widehat{S}_{+}$ and $\widehat{S}_{-}$ in equations Eq.(3),Eq.(5)the eigenfunctions of ”bra” and ”ket” states are obtained as
@@ V Conclusion and discussion
VI
 Acknowledgments
This work was supported by the National Natural Sci ence Foundation of China (Grant No. 12374312), ShanxiScholarship Council of China (Grant No. 2022-014 and 2023-033).
VII Conflict of Interest
The authors declare no conflict of interest.
VIII Data Availability Statement
The data that support the findings of this study are available from the cor responding author upon reasonable request.
IX Appendix
@@
Thus $\widehat{b}_{-}$ is a lowering operator for the “ket” state. We assume that
@@
Acting the rising operator on the Eq.(28) yields
@@
$c_{n-}^{r}$ $=$ $c_{\left(n-1\right)+}^{r}=\sqrt{n};c_{\left(n-1\right)-}^{r}=c_{\left(n-2\right)+}^{r}=\sqrt{n-1};\cdot\cdot\cdot;$
$c_{1-}^{r}$ $=$ $c_{0+}^{r}=1.$
Thus the $n$th “ket” state can be generated from ground state with the rising operator
@@
For the “bra” states situation is opposite that $\widehat{b}_{-}$ acts as a rising operator
@@
The recurrence relation for the “bra” states is obtained from the equations Eq.(8),Eq.(30) as
@@
$c_{n+}^{l}$ $=$ $c_{\left(n-1\right)-}^{l}=i\sqrt{n},c_{(n-1)+}^{l}=c_{\left(n-2\right)-}^{l}=i\sqrt{n-1};\cdot\cdot\cdot;$
$c_{1+}^{l}$ $=$ $c_{0-}^{l}=i,$
from which the $n$th “bra” state can be generated from the ground state such as
$|n\rangle_{l}=\frac{\left(\widehat{b}_{-}\right)^{n}}{i^{n}\sqrt{n!}}|0\rangle_{l}.$
@@
from which the ground “ket” state is found as
$\psi_{r,0}=\frac{1}{\sqrt{N}}e^{-i\frac{x^{2}}{2}},$
where $N$ is the normalization constant to be determined from the orthonormal condition Eq.(10) between “bra” and “ket” states. The arbitrary order eigenfunctions can be generated from the ground state such as
$\psi_{r,n}=\frac{1}{\sqrt{N}\sqrt{n!}}\left(\sqrt{\frac{i}{2}}\right)^{n}\left(x-i\frac{d}{dx}\right)^{n}e^{-i\frac{x^{2}}{2}}.$
With the same procedure we have the “bra” ground-state equation Eq.(31)
@@
$\psi_{l,0}=\frac{1}{\sqrt{N}}e^{i\frac{x^{2}}{2}}.$
The arbitrary-order wave function of “bra” state is
$\psi_{l,n}=\frac{1}{i^{n}\sqrt{n!}\sqrt{N}}\left(\sqrt{\frac{i}{2}}\right)^{n}\left(x+i\frac{d}{dx}\right)^{n}e^{i\frac{x^{2}}{2}}.$
The normalization constant can be determined from the orthonormal condition Eq.(10) between “bra” and “ket” states
@@
$N=\sqrt{\frac{\pi}{i}},$
is then obtained by means of the imaginary integral measure as in the path integral of quantum mechanics (see for example LW2023 ). The normalized wave functions are given by equation Eq.(12) in Sec.II.

### v2

@@ Abstract
April 3, 2025 April 3, 2025 April 3, 2025 April 3, 2025 April 3, 2025 April 3, 2025 April 3, 2025 April 3, 2025
@@ I Introduction
In quantum mechanics the Hamiltonian has to be Hermitian to guarantee real energy spectrum for the closed systems. The non-Hermitian Hamiltonian describes usually open systems, which exchange energy continuously with external source and thus possess complex energy spectrum. Bender and his collaborators proposed for the first time a creative finding that the Hermitian Hamiltonian is sufficient but not necessary to have real eigenvalues CM98 ; CM99 ; CM02 . According to them the non-Hermitian Hamiltonian with real spectrum has the parity-time ($PT$) reflection symmetry CM98 ; CM99 ; CM02 ; GuResult ; GuAnn ; LiuResult . Mostafazadeh named the non-Hermitian Hamiltonian with real spectrum as pseudo-Hermitian, which satisfies a necessary condition Mosta1 ; Mosta2 ; Mosta3 . While the $PT$ symmetry is neither necessary nor sufficient condition for a non-Hermitian Hamiltonian to have real spectrum. It is found that the $PT$asymmetric Hamiltonian can also possess real spectra GuResult ; GuAnn ; LiuPysica ; Amaa .
For the $PT$symmetric non-Hermitian Hamiltonians there exists an exceptional point, which separates the unbroken $PT$ symmetry region with pure real-spectrum and a broken $PT$ symmetry region, where eigenvalues are complex CM98 ; CM04 ; CM07 . This transition at the exceptional point has been observed in numerous laboratory experiments CM18 ; DN18 for various non-Hermitian Hamiltonians. The non-Hermitian Hamiltonian has become a rapidly developing subject in different branches of physicsCH13 ; ZZ16 ; SB12 ; CM13 ; CE10 ; YW19 The spectrum property becomes complicated at the exceptional pointsLC14 ; PG15 ; HX16 ; WL17 due to the symmetry breaking, which can be modulated by dissipative mediaSK19 ; JZ18 ; YZ17 ; FQ18 . The exceptional point and complex domain of eigenvalues have been well studied particularly in relation with optical systemsCM18 ; DN18 .
@@ II Hermitian Hamiltonian with an inverted potential well and imaginary spectrum
in which the subscripts ”$r$”, ”$l$” denote respectively the ”ket” and ”bra” states. The two sets of eigenstates are mutually orthonormal
@@
for both the ”bra” and ”ket” states ($\gamma=l,r$). The state densities $\langle x|\widehat{\rho}|x\rangle$, $\langle x|\widehat{\rho}^{{\dagger}}|x\rangle$ are conserved quantities satisfying the basic requirement of quantum mechanics.
@@
with $\gamma=r,l$ respectively for the ”ket” and ”bra” states. Indeed the Hermitian Hamiltonian with imaginary eigenvalues Eq.(11) has a dual-set of eigenstates in agreement with the general theory Eq.(1).
@@
$|n\rangle_{l}$ $=$ $\frac{\left(\widehat{b}_{-}\right)^{n}}{(-i)^{n}\sqrt{n!}}|0\rangle_{l}.$
@@
$\psi_{r,n}\left(x,\Omega\right)$ (12)
$=$ $\frac{1}{\sqrt{n!}}\left(\sqrt{\frac{i}{2}}\right)^{n}\left(\frac{i}{\pi}\right)^{\frac{1}{4}}\left(x-i\frac{d}{dx}\right)^{n}e^{-i\frac{x^{2}}{2}},$
$\psi_{l,n}\left(x,\Omega\right)$
$=$ $\frac{1}{\left(-i\right)^{n}\sqrt{n!}}\left(\sqrt{\frac{i}{2}}\right)^{n}\left(\frac{-i}{\pi}\right)^{\frac{1}{4}}\left(x+i\frac{d}{dx}\right)^{n}e^{i\frac{x^{2}}{2}}$
@@ III Transition from imaginary to real domains of spectrum and the exceptional point
where the real-value parameter $\eta$ is to be determined. The $SU(1,1)$ generators are realized by the imaginay-frequency boson operators in Eq.(5). It is easy to check that the transformation operator Eq.(17) is Hermitian $\widehat{R}=\widehat{R}^{{\dagger}}$ Using the transformation relations GuAnn LLM
$\widehat{R}\widehat{S}_{+}\widehat{R}^{-1}$ $=$ $\widehat{S}_{+}\cosh^{2}\frac{\eta}{2}+\widehat{S}_{-}\sinh^{2}\frac{\eta}{2}-i\widehat{S}_{z}\sinh\eta,$
$\widehat{R}\widehat{S}_{-}\widehat{R}^{-1}$ $=$ $\widehat{S}_{-}\cosh^{2}\frac{\eta}{2}-\widehat{S}_{+}\sinh^{2}\frac{\eta}{2}-i\widehat{S}_{z}\sinh\eta,$
$\widehat{R}\widehat{S}_{z}\widehat{R}^{-1}$ $=$ $\widehat{S}_{z}\cosh\eta+\frac{i}{2}(\widehat{S}_{-}-\widehat{S}_{+})\sinh\eta,$
@@
$\sinh\eta=\pm\frac{G}{\sqrt{\Omega^{2}-G^{2}}}$ (20)
@@
respectively for the ”ket” and ”bra” states, $\gamma=r,l.$ In our case it is simply the eigenstate of imaginary-frequency boson-number $\widehat{n}_{I}$
@@
defined in equation Eq.(9). The eigenstates of the original non-Hermitian Hamiltonians in equations Eq.(13,14) are obtained respectively by the inverse transforms
@@
Figure 2: The lowest layers ($n=0,1,2)$ of energy spectrum as functions of the interaction constant $G$. The imaginary eigenvalues below $G_{c}$ have two branches respectively for the ”ket” and ”bra” states and vanish at $G_{c}$. The spectrum becomes real beyond the exceptional point.
With the coordinate representation of $SU(1,1)$ generators $\widehat{S}_{+}$ and $\widehat{S}_{-}$ in equations Eqs.(3,5) the eigenfunctions of ”bra” and ”ket” states are obtained as
@@ V Conclusion and discussion
VI Appendix
@@
Thus $\widehat{b}_{-}$ is a lowering operator for the ”ket” state. We assume that
@@
Acting the raising operator on the Eq.(28) yields
@@
$c_{n-}^{r}=c_{\left(n-1\right)+}^{r}=\sqrt{n};c_{\left(n-1\right)-}^{r}=c_{\left(n-2\right)+}^{r}=\sqrt{n-1};\cdot\cdot\cdot;c_{1-}^{r}=c_{0+}^{r}=1.$
Thus the $n$th ”ket” state can be generated from ground state with the rising operator
@@
For the ”bra” states situation is opposite that $\widehat{b}_{-}$ acts as a raising operator
@@
The recurrence relation for the ”bra” states is obtained from the equations Eq.(8,30) as
@@
$c_{n+}^{l}$ $=$ $c_{\left(n-1\right)-}^{l}=-i\sqrt{n},c_{(n-1)+}^{l}=c_{\left(n-2\right)-}^{l}=-i\sqrt{n-1},$
$\cdot\cdot\cdot,c_{1+}^{l}$ $=$ $c_{0-}^{l}=-i,$
from which the $n$th ”bra” state can be generated from the ground state such as
$|n\rangle_{l}=\frac{\left(\widehat{b}_{-}\right)^{n}}{(-i)^{n}\sqrt{n!}}|0\rangle_{l}.$
@@
from which the ground ”ket” state is found as
$\psi_{r,0}=\frac{1}{\sqrt{N_{r}}}e^{-i\frac{x^{2}}{2}},$
where $N$ is the normalization constant to be determined from the orthonormal condition Eq.(10) between ”bra” and ”ket” states. The arbitrary order eigenfunctions can be generated from the ground state such as
$\psi_{r,n}=\frac{1}{\sqrt{N_{r}}\sqrt{n!}}\left(\sqrt{\frac{i}{2}}\right)^{n}\left(x-i\frac{d}{dx}\right)^{n}e^{-i\frac{x^{2}}{2}}.$
With the same procedure we have the ”bra” ground-state equation Eq.(31)
@@
$\psi_{l,0}=\frac{1}{\sqrt{N_{l}}}e^{i\frac{x^{2}}{2}}.$
The arbitrary-order wave function of ”bra” state is
$\psi_{l,n}=\frac{1}{(-i)^{n}\sqrt{n!}\sqrt{N_{l}}}\left(\sqrt{\frac{i}{2}}\right)^{n}\left(x+i\frac{d}{dx}\right)^{n}e^{i\frac{x^{2}}{2}}.$
The normalization constant can be determined from the orthonormal condition Eq.(10) between ”bra” and ”ket” states
@@
$N_{r}$ $=$ $\sqrt{\frac{\pi}{i}},$
$N_{l}$ $=$ $\sqrt{\frac{\pi}{-i}}$
is then obtained by means of the imaginary integral measure as in the path integral of quantum mechanics (see for example FH65 LW2023 ). The normalized wave functions are given by equation Eq.(12) in Sec.II.